[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sun May 28, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun May 28 18:31:06 UTC 2017


HSU Shift Log for Sun May 28 2017 14:30:59 GMT-0400 (EDT)

Shift: H

Forecaster: SR Stewart

Technical Issues Summary:

ATCF issues: None/None.

N-AWIPS issues: None/None.

AWIPS-2/NCP issues: None/None.

Other Techincal issues: None/None.

Miscellaneous issues: Performed pm on PC towers removing dust from fans and  
also cleaned out all keyboards.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170528/d9aefc63/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list